# Closing a catenary loop: the lariat chain, the string shooter, and the heavy elastica

### v2

@@ Abstract
We review, critique, and extend results related to the problem of closed loop shape equilibria of a string shooter, a type of catenary consisting of steady, axially moving configurations of an inertial, inextensible, perfectly flexible string in the presence of gravity and drag forces. We highlight recurring misconceptions, and relate to similar problems, including the lariat (no gravity), chain fountain (not closed), and heavy elastica (bending stiffness). We focus on the difficulty inherent to continuing a catenary through a vertical orientation, necessary to close a loop, which difficulty changes in nature as the system undergoes bifurcations with increasing drag. We construct solutions by implementing available analytical results, and numerically generate additional solutions with added bending stiffness. We briefly discuss global balances of linear, angular, and pseudo- momentum for this system.
@@ 1 Introduction
This extended comment, including a few new results, was stimulated by recent academic attention directed at a “cute” but challenging and historied problem in string mechanics, arising from its inclusion as part of problem 13 of the 11th International Physicists’ Tournament IPT19 , in turn inspired by a demonstration by Yeany YeanyYouTube . The problem is to determine the shape equilibria of a “string shooter”, a closed loop of axially moving string with a single support, subject to gravity and drag forces, to be described in detail below. As will be discussed at length, several important aspects of this and related problems, already documented throughout the literature, have been overlooked or misunderstood, and some recently proffered numerical solutions to the problem are not physically realizable. These conclusions follow from inspection of prior analytical results ChakrabartiHanna16 , however, errors in the initial presentation of these results must also be corrected in order to construct solutions. We will further connect with an earlier analytical result Gregory49 .
@@
We will provide a perspective on catenary problems that reviews both very old and somewhat new results, emphasizes a fundamental barrier to constructing solutions in the form of a closed loop with insufficient drag forces, and calls attention to the curious forms of singularities arising in this class of problems. Along the way, we must navigate past several distractions to reach the crux of the physics. In particular, we will show that the role of drag has been misrepresented, and perhaps over-emphasized, although indeed sufficient drag is necessary to construct loop solutions. Its role is to change tension and curvature profiles, at times through bifurcations, but it does not provide “lift”. On this point, classical results for axially moving bodies in the absence of drag, such as catenaries or free “lariats”, are illustrative.
@@
When drag forces are moderate to large, we indicate how analytical solutions can be used in a semi-analytical framework to construct closed loops. When drag forces are lower, such solutions are unphysical, and we offer a path towards numerical resolution of the problem using additional bending stiffness. This includes the problem of a static heavy elastica or, equivalently, a stiff catenary. Although bending stiffness, which is clearly present in some experiments, is a likely source of regularization, the nature of this regularization takes an atypical form. Instead of a high curvature boundary layer, derivatives of curvature contribute a bending force that allows the body to pass through an orientation aligned with gravity. We are unaware of any previous examination of this type of regularization. The reason we say “path towards” resolution is that the resulting problem of a very floppy beam is very stiff numerically, making the likely parameter range of experiments difficult to access. We do not attempt to match experiment to confirm that bending is indeed the regularizing agent in the laboratory, but show that a small amount of bending has a large qualitative effect on the solutions.
@@ 2 Overview of the problem and its solutions
Briefly, the problem at hand considers observed steady or nearly-steady configurations of an axially moving loop of string or chain subject to gravitational and drag forces. Most of the body is suspended in air, but at one location, one or more motorized wheels drive the axial motion, as well as provide a necessary force to support the structure. An example of such a flowing structure is shown in Figure 1. For simplicity, we may presume perfect flexibility, inextensibility, and uniform density of a continuous body. What, if any, are the equilibrium loop shapes and corresponding tensions, and how do they depend on drag forces? Here by “loop” we mean a closed curve that might have kinks (jumps in angle), particularly at the support point. The tangents of such a shape must necessarily either subtend an angle greater than $\pi$, including a vertical point, or have kinks. However, jumps in internal force and moment are only admissible at the support point, where external supplies are available. We anticipate that the supply of force will be vertical (opposing gravity), and expect that the supply of moment should be zero for a perfectly flexible string.
Figure 1: A string shooter in operation, and its inventor StringLauncher . The string moves axially clockwise with speed $v$. The launch mechanism provides a force $\bf{R}$}. (Annotations added to a still from YeanyYouTube , flipped horizontally for comparison with other figures in the present paper.)
The analytical results of ChakrabartiHanna16 can be used to generate solutions for sufficiently high drag forces. For lower drag forces, there are no solutions to the problem as posed above.
To understand why, it is helpful to review the classical case of a static catenary, in which tension balances gravity. The string’s contribution to normal force balance is the product of tension and curvature. There are two consequences of this: any nonvertical element of string cannot be straight or tensionless, and any vertical element of string must either be straight or tensionless or both, otherwise there will be an unbalanced force. With the exception of nongeneric vertical straight line solutions, a hanging catenary approaches the vertical asymptotically, with tension and arc length tending to infinity and curvature tending to zero. The same is true for a catenary arch, except that the tension is of negative sign. Centripetal acceleration associated with axial motion will shift the tension by some positive constant, which can affect stability, but does not change the shape.
@@
However, stronger drag forces allow for physical patching, through two different scenarios, depending on the magnitude of the drag. Three regimes, separated by two bifurcations noted but misinterpreted in ChakrabartiHanna16 , are
@@
The results of patching will be seen in greater detail below. The symmetry or asymmetry of the resulting shapes depends on the interrelated launch and return angles of the shot string, but it is a curious property of the catenary equations that the lengths of the upper and lower curve branches will be equal, even for asymmetric solutions Gregory49 . A given launch angle and string length will determine a particular spatial location for the patch point, and a particular return angle.
@@ 3 A bit of history
Aitken’s analysis of his experiments benefited from prior knowledge of relevant mechanics. In his own modest assessment, “The experiments do not contain much that is new, many of the problems having previously been mathematically wrought out by Sir William Thomson and others” Aitken1878 . The dominant normal force balance between centripetal acceleration and tension was well understood in the 1850s by Routh Routh55 , who may have picked this up from the Mathematical Tripos. Both of these contributions depend linearly on curvature, which can be divided out of the normal force balance, such that no information about shape is retained and any configuration of the axially moving body is a marginally stable equilibrium. One can alternately recognize that one of two arbitrarily shaped nonlinear wave solutions of an axially moving string is “frozen” in the lab frame. The only effect of axial motion is to positively shift the value of internal tension. This is generally true for static shape equilibria in which no drag or Coriolis forces are present. In particular, catenaries with axial flow retain their shape, with tension shifted, as noted by Routh in his treatise Routh55 , and in contemporary analyses of telegraph-laying by Airy Airy1858 and Thomson Thomson1857both ; for a review of the literature on paying out and towing cables with additional drag forces, see ChakrabartiHanna16 . For more on shape-agnosticism, tension shifts, and other interesting discussions, we refer the reader to Routh [Routh55, , Chapter XIII, Articles 594-598], Gray [Gray59, , Chapter XXII, Sections 11-13], and a short note by Healey and Papadopoulos that extends this “classical, but virtually forgotten, result” HealeyPapadopoulos90 . A striking consequence of the shift in tension due to axial flow is the stabilization of structures lacking bending stiffness that would otherwise collapse if under compression, in particular the catenary arch. More than a century after Aitken’s demonstrations, this phenomenon was rediscovered, explained, and experimentally demonstrated by Perkins and Mote PerkinsMote89 and again later by others investigating the “chain fountain” or closely related phenomena HannaKing11 , HannaSantangelo12 , Biggins14 .
For lighter bodies and higher velocities, which have higher drag to weight ratios, drag plays a more noticeable role, affecting both shape and tension. Slow shape dynamics and wave behavior are still interesting features, constitute the remainder of the IPT problem IPT19 , and are on full display in Yeany’s “String Shooter/Launcher” toy and science demonstration StringLauncher , YeanyYouTube . An analytical solution, accompanied by some experiments, was put forward by the astronomer Gregory in the 1940s Gregory49 .222…not long before retiring from science to pursue interests in parapsychology GregoryObit1 , GregoryObit2 . Gregory cites no previous work— later authors will reciprocate— but indicates the following origin: “It has long been known to engineers that a belt when driven at a high speed will occasionally be seen to stand up in the air in the form of a loop… The author’s attention was drawn to this problem at the Admiralty when examining a proposal to use such a loop as a means of defence against aircraft.” He expresses interest in using this problem to measure drag forces. A similar sentiment, connected with the physics of fly fishing lines, drove the recent lariat chain experiments of Judd JuddLariat . Gregory patches upper and lower curves together to create a loop. He realizes that at lower drag the tension is clearly not patchable, and that at higher drag, the curvature becomes singular at the patch point. He thus interprets his analytical solution as being only valid for intermediate values of drag, not recognizing the potential validity of singular high drag solutions.
@@ 4 Critique of recent literature
Most of what is needed to understand the string shooter problem is present in ChakrabartiHanna16 , as a special case of a broader multiparameter family of solutions that include translations of the catenary at any angle with respect to gravity. While anisotropic drag linear in velocity was considered, it was noted that for purely axially moving bodies, the solutions are more generally valid, as the drag is tangential and uniform regardless of its dependence on velocity. The solutions undergo two important bifurcations in tension and shape, whose implications for patching were given above in Section 2. The first of these, occurring at a drag to weight ratio of unity, allows the inertially-shifted tension to vanish at finite length on one side. The second of these, occurring at twice this ratio, leads to finite arc length curvature blowup. The authors of ChakrabartiHanna16 , including one of the authors of the present paper, made two errors. First, and most importantly, they did not realize that the first bifurcation leads to an integrable singularity, allowing patching at finite length. This led to strange results involving finite tensions despite infinite lengths, which are likely incorrect and require revisiting. [^3] Contributing to this confusion were obstacles to continuing numerical solutions through vertical orientation, and the conceptualization of the solutions in terms of heteroclinic connections in “phase portraits” for angle and curvature; above the first bifurcation, the solutions reach the angle axis orthogonally instead of asymptotically approaching it, and are thus no longer unique there, but the corresponding change in appearance of the portrait is subtle. This error in interpreting the first bifurcation carried over into an aspect of the second, where it seemed that a special nonsingular solution existed right at the second bifurcation. As stated in ChakrabartiHanna16 , “The existence of such a solution only under such restrictive conditions on speed is surprising, and this result will remain suspect until it is observed experimentally.” This is now resolved by the recognition that the suspect result was simply wrong. Second, they made an incorrect statement to the effect that the change in tension from infinite to zero occurs at the leading rather than the trailing side of the catenary with respect to the axial flow [ChakrabartiHanna16, , (29)], when in fact the location of this change depends on the parameters of the problem; were the statement strictly correct, we would be unable to patch one leading upper and one trailing lower side as is necessary here to create our loops. Note that transitions at both sides are possible for the broader multiparameter family, but not for the present problem, which sits along the right edge of the diagram in [ChakrabartiHanna16, , figure 13].
@@
Bevan and Deane BevanDeane19 rediscover the solution from ChakrabartiHanna16 and use it to describe the “chain fountain”, a catenary arch with tangential drag forces. They don’t realize that the drag force need not be linear. They do realize that low drag solutions cannot reach vertical orientation, which they do not need for their problem, and note the significance of the first bifurcation, but do not admit solutions above the second. Abello and co-workers Abello20 find what appears to be a Cartesian parameterization of the solution, which they interpret as being only valid above the first bifurcation.
@@
We conclude this section by addressing the relevant portions of two papers from 2019 that provide numerical solutions, and our interpretation of these in the light of previously published analytical solutions and bifurcations Gregory49 , ChakrabartiHanna16 which we believe clarify some of the observations. The two groups appear to be unaware of available analytical solutions and their interpretations, including the bifurcation generating the high drag curvature singularity.
@@
Taberlet and co-workers Taberlet19 consider two paradigms, guided by experiments on heavy beaded chains and light strings draped over a pulley. For heavy bodies, they ignore drag, and further ignore the difficulties associated with vertical orientation by assuming such points sit on the pulley. They mistakenly say that the shifted tension at the bottom of the catenary is zero, which leaves nothing to balance gravity. For light bodies they ignore gravity— the source of difficulty— so that verticality is no longer forbidden. However, the only steady shapes with drag and without gravity are straight lines ChakrabartiHanna16 . For general cases between these limits, numerical solutions to the steady-state balance equations are generated using an unspecified method to “avoid the singularity”. A shape is shown for high drag, and data is presented for both moderate and high drag. They also perform dynamic numerical simulations using 1000 beads and springs, for low through high drag values. They infer the presence of the first bifurcation from the behavior of their solutions, but seem to be unaware of the second bifurcation, above which most of their results lie. It is not clear whether the low drag simulations reach steady state, or whether an effective “bending” due to discretization, or some dynamic effect, can allow closed loop solutions below the first bifurcation. They report a numerical balance between the moments due to drag and weight forces, which is as expected if no moment is supplied at the support, and there is no net angular acceleration of the loop. The effects of drag and launch angle on shape and orientation are shown. An experimental image in their supplemental material shows a “dolphin nose” shape Judddolphin which, along with other features visible throughout, is presumably due to bending stiffness.
@@
Daerr and co-workers Daerr19 numerically obtain a family of static string solutions for upper and lower parts of a loop for moderate and high drag, which they patch, guided by experimental shapes. Asymptotic behavior near the vertical is considered analytically; this expansion would not be possible at low drag due to the divergence of arc length at the vertical. Although bifurcations can be inferred from these expressions, they seem unaware of them, although they note qualitative differences in shape at high drag. They employ two approaches. In one, they patch upper and lower curves at the vertical, empirically rediscovering Gregory’s result of equal lengths. In the other, they create a large jump in angle by cutting off the small region of high curvature in the top part of the experimental curves (see e.g. their figure 5c). They interpret this kink as something physical and measurable, saying the solution is obtained “on condition that both the angle and the tension are allowed to be discontinuous”, but this is not physically admissible, as it requires an injection of force from some external source not present. We interpret the “apparent angle discontinuity” data reported in their figure 5a as an arbitrary choice of these authors, rather than a physical effect. They later invoke the possibility of nonlocal drag forces as the regularizer of this feature.

### v3

@@ Abstract
We review, critique, and extend results related to the problem of closed loop shape equilibria of a string shooter, a type of catenary consisting of steady, axially moving configurations of an inertial, inextensible, perfectly flexible string in the presence of gravity and drag forces. We relate to similar problems, including the lariat (no gravity), chain fountain (not closed), and heavy elastica (bending stiffness). We focus on the difficulty inherent to continuing a catenary through a vertical orientation, necessary to close a loop, which difficulty changes in nature as the system undergoes bifurcations with increasing drag. We construct solutions by implementing available analytical results, and numerically generate additional solutions with added bending stiffness. We briefly discuss global balances of linear, angular, and pseudo- momentum for this system.
@@ 1 Introduction
This extended comment, including a few new results, was stimulated by recent academic attention directed at a “cute” but challenging and historied problem in string mechanics, arising from its inclusion as part of problem 13 of the 11th International Physicists’ Tournament IPT19 , in turn inspired by a demonstration by Yeany YeanyYouTube . The problem is to determine the shape equilibria of a “string shooter”, a closed loop of axially moving string with a single support, subject to gravity and drag forces, to be described in detail below. As will be discussed at length, several important aspects of this and related problems, already documented throughout the literature, have been overlooked, and some recently published numerical solutions to the problem are not physically realizable. These conclusions follow from inspection of prior analytical results Gregory49 , SvetlitskiiMiroshnik73 , Miroshnik01 , ChakrabartiHanna16 , however, errors in the presentation of the results in ChakrabartiHanna16 must also be corrected in order to construct solutions.
@@
We will provide a perspective on catenary problems that reviews both very old and somewhat new results, emphasizes a fundamental barrier to constructing solutions in the form of a closed loop with insufficient drag forces, and calls attention to the curious forms of singularities arising in this class of problems. Along the way, we must navigate past several distractions to reach the crux of the physics. In particular, from our perspective, the importance of drag has been misunderstood, and perhaps over-emphasized. Although sufficient drag is necessary to construct loop solutions, its role is to change tension and curvature profiles, at times through bifurcations. On this point, classical results for axially moving bodies in the absence of drag, such as catenaries or free “lariats”, are illustrative.
@@
When drag forces are moderate to large, we indicate how analytical solutions can be used in a semi-analytical framework to construct closed loops. When drag forces are lower, such solutions are unphysical, and we offer a path towards numerical resolution of the problem using additional bending stiffness. This includes the problem of a static heavy elastica or, equivalently, a stiff catenary. Although bending stiffness, which is clearly present in some experiments, is a likely source of regularization, the nature of this regularization takes an atypical form. Instead of a high curvature boundary layer, derivatives of curvature contribute a bending force that allows the body to pass through an orientation aligned with gravity. We are unaware of any previous examination of this type of regularization. The reason we say “path towards” resolution is that the resulting problem of a very floppy beam is very stiff numerically, making the likely parameter range of experiments difficult to access. We show that a small amount of bending has a large qualitative effect on the solutions. We do not attempt to match experiment to determine whether bending is indeed the regularizing agent in the laboratory.
@@ 2 Overview of the problem and its solutions
Briefly, the problem at hand considers observed steady or nearly-steady configurations of an axially moving loop of string or chain subject to gravity and drag forces. Most of the body is suspended in air, but at one location, one or more motorized wheels drive the axial motion, as well as provide a necessary force to support the structure. An example of such a flowing structure is shown in Figure 1. For simplicity, we may presume perfect flexibility, inextensibility, and uniform density of a continuous body, coplanar with gravity. What, if any, are the equilibrium loop shapes and corresponding tensions, and how do they depend on drag forces? Here by “loop” we mean a closed curve that might have kinks (jumps in angle), particularly at the support point. The tangents of such a shape must necessarily either have two or more kinks, or subtend an angle greater than $\pi$, including a vertical point (where the normal is horizontal). However, jumps in internal force and moment are only admissible at the support point, where external supplies are available. We anticipate that the supply of force will be vertical (opposing gravity), and expect that the supply of moment should be zero for a perfectly flexible string.
Figure 1: A string shooter in operation, and its inventor StringLauncher . The string moves axially clockwise with speed $v$. The launch mechanism provides a force $\bf{R}$}. (Annotations added to a still from YeanyYouTube , flipped horizontally for consistency with other figures in the present paper.)
The analytical results of Gregory49 , SvetlitskiiMiroshnik73 , Miroshnik01 , ChakrabartiHanna16 can be used to generate solutions for sufficiently high drag forces. For lower drag forces, there are no solutions to the problem as posed above.
To understand why, it is helpful to review the classical case of a static catenary, in which tension balances gravity. The string’s contribution to normal force balance is the product of tension and curvature. There are two consequences of this: any nonvertical element of string cannot be straight or tensionless, and any vertical element of string must either be straight or tensionless or both, otherwise there will be an unbalanced force. With the exception of nongeneric vertical straight line solutions, a hanging catenary approaches the vertical asymptotically, with tension and arc length tending to infinity, and curvature tending to zero. The same is true for a catenary arch, with the tension negative. Centripetal acceleration associated with axial motion will shift the tension by some positive constant, which can affect stability, but does not change the shape.
@@
However, stronger drag forces allow for physical patching, through two different scenarios, depending on the magnitude of the drag. As noted in SvetlitskiiMiroshnik73 , Miroshnik01 , ChakrabartiHanna16 , although misinterpreted in ChakrabartiHanna16 , there are three regimes, separated by two bifurcations:
@@
The results of patching will be seen in greater detail below. The symmetry or asymmetry of the shapes depends on the interrelated launch and return angles of the shot string, but it is a curious property of the catenary equations that the lengths of the upper and lower curve branches will be equal, even for asymmetric solutions Gregory49 . A given launch angle and string length will determine a particular spatial location for the patch point, and a particular return angle.
@@ 3 A bit of history
Aitken’s analysis of his experiments benefited from prior knowledge of relevant mechanics. In his own modest assessment, “The experiments do not contain much that is new, many of the problems having previously been mathematically wrought out by Sir William Thomson and others” Aitken1878 . The dominant normal force balance between centripetal acceleration and tension was well understood in the 1850s by Routh Routh55 , who may have picked this up from the Mathematical Tripos. Both of these contributions depend linearly on curvature, which can be divided out of the normal force balance, such that no information about shape is retained and any configuration of the axially moving body is a marginally stable equilibrium. One can alternately recognize that one of two arbitrarily shaped nonlinear wave solutions of an axially moving string is “frozen” in the lab frame. The only effect of axial motion is to positively shift the value of internal tension. This is generally true for static shape equilibria in which no drag or Coriolis forces are present. In particular, catenaries with axial flow retain their shape, with tension shifted, as noted by Routh in his treatise Routh55 , and in contemporary analyses of telegraph-laying by Airy Airy1858 and Thomson Thomson1857both ; for a review of the literature on paying out and towing cables with additional drag forces, see ChakrabartiHanna16 and the brief mention of normal drag in the paragraph below. For more on shape-agnosticism, tension shifts, and other interesting discussions, we refer the reader to Routh [Routh55, , Chapter XIII, Articles 594-598], Gray [Gray59, , Chapter XXII, Sections 11-13], and a short note by Healey and Papadopoulos that extends this “classical, but virtually forgotten, result” HealeyPapadopoulos90 . A striking consequence of the shift in tension due to axial flow is the stabilization of structures lacking bending stiffness that would otherwise collapse if under compression, in particular the catenary arch. More than a century after Aitken’s demonstrations, this phenomenon was rediscovered, explained, and experimentally demonstrated by Perkins and Mote PerkinsMote89 and again later by others investigating the “chain fountain” or closely related phenomena HannaKing11 , HannaSantangelo12 , Biggins14 .
For lighter bodies and higher velocities, which have higher drag to weight ratios, drag plays a more noticeable role, affecting both shape and tension. Slow shape dynamics and wave behavior are still interesting features, constitute the remainder of the IPT problem IPT19 , and are on full display in Yeany’s “String Shooter/Launcher” toy and science demonstration StringLauncher , YeanyYouTube . An analytical solution for curvature and tension in terms of tangential angle was put forward by the astronomer Gregory in the 1940s Gregory49 , along with a brief report of some experiments.222…not long before retiring from science to pursue interests in parapsychology GregoryObit1 , GregoryObit2 . Gregory cites no previous work— later authors will reciprocate— but indicates the following origin: “It has long been known to engineers that a belt when driven at a high speed will occasionally be seen to stand up in the air in the form of a loop… The author’s attention was drawn to this problem at the Admiralty when examining a proposal to use such a loop as a means of defence against aircraft.” He expresses interest in using this problem to measure drag forces. A similar sentiment, connected with the physics of fly fishing lines, drove the recent lariat chain experiments of Judd JuddLariat . Gregory patches upper and lower curves together to create a loop. He realizes that at lower drag the tension is clearly not patchable, and that at higher drag, the curvature becomes singular at the patch point. He thus interprets his analytical solution as being only valid for intermediate values of drag, not recognizing the potential validity of singular high drag solutions. Analytical and experimental results appear in an independent body of work beginning in the 1960s Kurkin64 , SvetlitskiiGabryuk66 , Svetlitskii72inrussian , Svetlitskii72 , Miroshnik72inrussian , SvetlitskiiMiroshnik73 , MiroshnikKurkin76inrussian , Miroshnik01 , [^3] This body of work has gone uncited by later investigators of the string shooter, and the present authors only became aware of it during late-stage revisions. These works also cite earlier and concurrent theoretical and experimental results by other authors, in references which we were unable to obtain at the time of writing. inspired by antenna and radiator applications, and at times treating additional complications such as a normal drag force quadratic in velocity, a fluid-fluid interface, or a finite distance between launch and return points. Kurkin Kurkin64 expresses tension in terms of tangential angle, with arc length and thus any shape information remaining in quadrature. He reports experiments in vacuum demonstrating that increasing drag leads to narrower and higher loops, but does not explicitly recognize any transitions. He proposes measuring the drag coefficient by using the balance of moments about the support. Svetlitskii and Gabryuk SvetlitskiiGabryuk66 find an implicit solution relating Cartesian coordinates and arc length, and show experimental shapes for moderate and high drag, noting that low drag solutions are not admissible. These shapes display features suggestive of bending stiffness. As they do not solve for the curvature, they do not notice the second bifurcation. No theoretical curves are shown, and patching at the vertical is not discussed. Svetlitskii, Miroshnik, and Kurkin Svetlitskii72inrussian , Svetlitskii72 find an explicit Cartesian parameterization and expression for the tension, and favorably compare a high drag experimental shape with a solution, estimating the drag coefficient with Kurkin’s method. They remain unaware of the second bifurcation, and do not provide a clear discussion of patching. Miroshnik Miroshnik72inrussian expresses tension in terms of tangential angle. The string shooter is a special case of his more general solution, which features an additional purely normal drag force quadratic in velocity. He notes a singularity at the vertical and identifies the first bifurcation, showing high drag solutions while apparently unaware of the second bifurcation. The effects of angle and drag are shown. Examples include a vertically launched shape, which is impossible in the absence of normal drag— it’s not clear whether the author recognized this and verified the solution, but this point was made in Svetlitskii72inrussian . Continuing work on this general solution, Svetlitskii and Miroshnik SvetlitskiiMiroshnik73 express curvature and tension in terms of tangential angle, thereby identifying both bifurcations, the corresponding behaviors at the vertical point, and the difference between moderate and high drag shapes. Miroshnik and Kurkin MiroshnikKurkin76inrussian make use of known Cartesian solution forms for parameterization and tension and add to these an expression for arc length. By considering independent launch and “return” points, beyond simply two locations on a pulley, they explore tent-shaped and S-shaped solutions, and shape transitions under changes in drag and string length. Decades later, Miroshnik Miroshnik01 revisits and discusses many aspects of the problem, including the bifurcations and the vertical singularity, for similarly general boundary conditions. This important, yet overlooked, paper provides a Cartesian parameterization and expressions for tension, arc length, and curvature, and favorably compares experimental shapes with solutions for different drag regimes and launch angles, again estimating drag coefficients with Kurkin’s method. The experimental shapes display features suggestive of bending stiffness, including a “dolphin nose” Judddolphin in his figure 10. Hanging low drag catenaries are also discussed. Some examples of tension as a function of arc length are (parametrically) plotted. Miroshnik also suggests two additional “modes” for which the string’s direction of motion is ascending through a vertical singularity, and dismisses them because they self-intersect. However, such an ascending singularity is not possible because the arc length and tension are not regularized there, something this paper does not recognize, despite the observation in Svetlitskii72inrussian that a vertical launch is not possible.
@@ 4 Critique of recent literature
Most of what is needed to understand the abstract string shooter problem is present in Gregory49 , SvetlitskiiMiroshnik73 , Miroshnik01 , ChakrabartiHanna16 . In ChakrabartiHanna16 , the problem appears as a special case of a broader multiparameter family of solutions that include translations of the catenary at any angle with respect to gravity. While anisotropic drag linear in velocity was considered, it was noted that for purely axially moving bodies, the solutions are more generally valid, as the drag is tangential and uniform regardless of its dependence on velocity. The solutions undergo two important bifurcations in tension and shape, whose implications for patching were given above in Section 2. The first of these, occurring at a drag to weight ratio of unity, allows the inertially-shifted tension to vanish at finite length on one side. The second of these, occurring at twice this ratio, leads to finite arc length curvature blowup. However, the authors of ChakrabartiHanna16 , including one of the authors of the present paper, made two errors. First, and most importantly, they did not realize that the first bifurcation leads to an integrable singularity, allowing patching at finite length. This led to strange results involving finite tensions despite infinite lengths, which are likely incorrect and require revisiting. [^4] Contributing to this confusion were obstacles to continuing numerical solutions through vertical orientation, and the conceptualization of the solutions in terms of heteroclinic connections in “phase portraits” for angle and curvature; above the first bifurcation, the solutions reach the angle axis orthogonally instead of asymptotically approaching it, and are thus no longer unique there, but the corresponding change in appearance of the portrait is subtle. This error in interpreting the first bifurcation carried over into an aspect of the second, where it seemed that a special nonsingular solution existed right at the second bifurcation. As stated in ChakrabartiHanna16 , “The existence of such a solution only under such restrictive conditions on speed is surprising, and this result will remain suspect until it is observed experimentally.” This is now resolved by the recognition that the suspect result was simply wrong. Second, they made an incorrect statement to the effect that the change in tension from infinite to zero occurs at the leading rather than the trailing side of the catenary with respect to the axial flow [ChakrabartiHanna16, , (29)], when in fact the location of this change depends on the parameters of the problem; were the statement strictly correct, we would be unable to patch one leading upper and one trailing lower side as is necessary here to create our loops. Note that transitions at both sides are possible for the broader multiparameter family, but not for the present problem, which sits along the right edge of the diagram in [ChakrabartiHanna16, , figure 13]. This asymmetry disallows the additional modes suggested in Miroshnik01 .
@@
Bevan and Deane BevanDeane19 rediscover the solution and use it to describe the “chain fountain”, a catenary arch with tangential drag forces. They don’t realize that the drag force need not be linear. They do realize that low drag solutions cannot reach vertical orientation, which they do not need for their problem. They do note the significance of the first bifurcation, but, like Gregory before them, do not admit solutions above the second. Abello and co-workers Abello20 rediscover the Cartesian parameterization, which they interpret as being only valid above the first bifurcation.
@@
We conclude this section by addressing the relevant portions of two papers from 2019 that provide numerical solutions, and our interpretation of these in the light of previously published analytical solutions and bifurcations Gregory49 , SvetlitskiiMiroshnik73 , Miroshnik01 , ChakrabartiHanna16 which we believe clarify some of the observations. The two groups appear to be unaware of available analytical solutions and their interpretations, including the bifurcation generating the high drag curvature singularity.
@@
Taberlet and co-workers Taberlet19 consider two paradigms, guided by experiments on heavy beaded chains and light strings draped over a pulley. For heavy bodies, they ignore drag, and further ignore the difficulties associated with vertical orientation by assuming such points sit on the pulley. They mistakenly say that the shifted tension at the bottom of the catenary is zero, which leaves nothing to balance gravity. For light bodies they ignore gravity— the source of difficulty— so that verticality is no longer forbidden. However, the only steady shapes with drag and without gravity are straight lines Svetlitskii72 , ChakrabartiHanna16 . For general cases between these limits, numerical solutions to the steady-state balance equations are generated using an unspecified method to “avoid the singularity”. A shape is shown for high drag, and data is presented for both moderate and high drag. They also perform dynamic numerical simulations using 1000 beads and springs, for low through high drag values. They infer the presence of the first bifurcation from the behavior of their solutions, but seem to be unaware of the second bifurcation, above which most of their results lie. It is not clear whether the low drag simulations reach steady state, or whether an effective “bending” due to discretization, or some dynamic effect, can allow closed loop solutions below the first bifurcation. They report a numerical balance between the moments due to drag and weight forces, which is as expected if no moment is supplied at the support, and there is no net angular acceleration of the loop. The effects of drag and launch angle on shape and orientation are shown. An experimental image in their supplemental material shows a “dolphin nose” shape Judddolphin which, along with other features visible throughout, is presumably due to bending stiffness.
@@
Daerr and co-workers Daerr19 numerically obtain a family of static string solutions for upper and lower parts of a loop for moderate and high drag, which they patch, guided by experimental shapes. Asymptotic behavior near the vertical is considered analytically; this expansion would not be possible at low drag due to the divergence of arc length at the vertical. Although bifurcations can be inferred from these expressions, they seem unaware of them, although they note qualitative differences in shape at high drag. They rediscover Kurkin’s method for measuring drag coefficients. They employ two approaches to form loops. In one, they patch upper and lower curves at the vertical, empirically rediscovering Gregory’s result of equal lengths. In the other, they create a large jump in angle by cutting off the small region of high curvature in the top part of the experimental curves (see e.g. their figure 5c). They interpret this kink as something physical and measurable, saying the solution is obtained “on condition that both the angle and the tension are allowed to be discontinuous”, but this is not physically admissible, as it requires an injection of force from some external source not present. We interpret the “apparent angle discontinuity” data reported in their figure 5a as an arbitrary choice of these authors, rather than a physical effect. They later invoke the possibility of nonlocal drag forces as the regularizer of this feature.